Calculate Payroll 

An employee earns $30 per hour and 1.5 times that rate for all hours in excess of 40 hours per week. If the employee worked 54 hours during the week. Assume that the social security tax rate is 6.0%, the Medicare tax rate is 1.5%, and the employee's federal income tax withheld is $295.

a. Determine the employee's gross pay for the week. 

b. Determine the employee's net pay for the week. Round your answer to two decimal places.
Answer #2

Correct Answer:

Requirement 1:

Gross Pay = $ 1,830.00

Working:

pay per hour *

No. of hours worked in a week =

Total Pay for the week

30

40

$ 1200

45

14

$ 630

Gross Pay

$ 1830

Requirement 2:

Net pay = $ 1,397.75

Working:

Gross Pay

$    1,830.00

Less: Federal Income Tax

$        295.00

Less: Social Security @ 6.0% of 1830

$        109.80

Less: Medicare tax @ 1.5% of 1830

$           27.45

Total Tax deductible

$        432.25

Net pay

$    1,397.75

End of answer.

Please give thumbs up, it will be highly appreciated.

Thanks.
Answer #1

Part A

Gross pay = regular pay + overtime pay = (40*30)+(14*30*1.5) = $1830

Part B

Net pay = gross pay – social security tax – medicare tax – federal income tax = 1830-(1830*6.0%)-(1830*1.5%)-295 = $1397.75
Answer #3

a. Gross pay for the week =(40*30)+(14*30*1.5) = $1830

b. Net pay = $1830 - 295 - (1830*7.5%) = $1400
Answer #4

A)

Gross pay = (40*30) + (14*30*1.50)

= 1830

B)

Net pay = Gross pay - Social security tax - Medicare tax - federal income tax withheld

= 1830 - (1830*6%) - (1830*1.5%) - 295

= 1397.75